OpenStudy (anonymous):

If g(x)+x sin g(x)=x^2 and g(1)=0, find g′(1).

OpenStudy (zugzwang):

g(1) = 0 ?? Then g(1) + x sin g(1) = 1^2 0 + 1 sin 0 = 1 0 = 1 Am I missing something here?

OpenStudy (anonymous):

im looking for the derivative of g(1)

OpenStudy (anonymous):

lol how i wished that was the question

OpenStudy (shubhamsrg):

what @zugzwang means is that probably there's a typo or ques is incorrect..

OpenStudy (anonymous):

ahhh i see ya sorry i did type the question out incorrectly let me fix that